EXPERIMENTAL AND THEORETICAL STUDIES OF ETHYL

\underline{TREVOR J. SEARS}, Department of Chemistry, Brookhaven National Laboratory, Upton, NY 11973; and PHILIP M. JOHNSON, Department of Chemistry, SUNY, Stony Brook, NY 11794.

First detected in 1988, the −CH_{2} wagging fundamental of C_{2}H_{5} recorded at high resolution by diode laser absorption extends from approximately 470 to more than 600 cm^{−1}. The mode corresponds to the out-of-plane motion at the radical center and this is strongly coupled to the torsion or internal rotation in this species. We published^{a} a preliminary analysis of the spectrum in 1996, but at the time recognized many questions remained. Aided by extensive ab initio calculations and the detection^{b} of part of the spectrum at 3μm in a jet-cooled sample, we have recently made much progress understanding the details of the 20μm spectrum and the internal dynamics in the radical. In total, close to 500 transitions have been assigned in 18 rotation-torsion branches. The observation of branches involving levels with m_{torsion} = 3 and low k_{a}, in particular, has allowed much more precise estimates of the barrier to internal rotation and its change on vibrational excitation. Calculations of the torsion/wag potential surface have allowed a physical interpretation of these experimental results. Acknowledgement: This work was carried out at Brookhaven National Laboratory under contract DE-AC02-98CH10886 with the U. S. Department of Energy and supported by its Division of Chemical Sciences, Office of Basic Energy Sciences.

aT. J. Sears, P. M. Johnson, P. Jin and S. Oatis, J. Chem. Phys., 104, 781 (1996)

bS. Davis, D. Uy, and D. J. Nesbitt, Paper FE03, 52^{nd} Ohio State University International Symposium on Molecular Spectroscopy,